3b) Calculate the noise power and noise voltage generated by a 1M resistor at 37C 

over a 1 MHz bandwidth.


1
Expert's answer
2021-03-29T09:06:25-0400

The noise power is given by the following formula (see https://en.wikipedia.org/wiki/Johnson–Nyquist_noise):



"P =k_BT\Delta f"

where "k_B = 1.38\times 10^{-23}J/K" is the Boltzmann's constant, "T = 37\degree C = 310K" is the absolute temperature, and "\Delta f = 1\times 10^{6}Hz" is the bandwidth. Thus, obtian:



"P = 1.38\times 10^{-23}\cdot 310K\cdot 10^{6} = 4.28\times 10^{-15}W"

The noise voltage is:



"E = \sqrt{4Rk_BT\Delta f}"

where "R = 1M\Omega = 10^{6}\Omega" is the resistance. Thus, obtain:



"E = \sqrt{4\cdot 10^6\cdot 1.38\times 10^{-23}\cdot 310\cdot 10^6} \approx 1.31\times 10^{-4}V"

Answer. "4.28\times 10^{-15}W" and "1.31\times 10^{-4}V".